# A bound-preserving and conservative enriched Galerkin method for elliptic problems

### v1

@@ Abstract
We propose a locally conservative enriched Galerkin scheme that respects the discrete maximum principle of an elliptic problem. To this end, we use a substantial over-penalization of the discrete solution’s jumps to obtain optimal convergence. To avoid the ill-conditioning issues that arise in over-penalized schemes, we introduce an involved splitting approach that separates the system of equations for the discontinuous solution part from the system of equations for the continuous solution part, yielding well-behaved subproblems. We prove the existence of discrete solutions and optimal error estimates, which are validated numerically.
@@ 1 Introduction
Physically-consistent numerical methods have become a key focus in the numerical analysis of partial differential equations (PDEs). These methods aim to preserve specific properties of the exact solutions, such as local conservation laws, entropy conditions, discrete maximum principles, or positivity constraints. Such properties are critical for ensuring that numerical simulations of complex phenomena remain physically accurate and stable.
In particular, methods that satisfy the Discrete Maximum Principle (DMP) and monotonicity conditions have been extensively studied within the finite element framework [Cia70, CR73, Kik77, MH85, BE05, BKK08, BBK17] (see also [BJK25] for a recent review on the topic). These properties, and the more general bound preservation criterion, guarantee that numerical solutions adhere to the inherent bounds of the exact solution, often preserving positivity and/or other relevant physical bounds. Ensuring that these numerical methods respect such constraints is crucial, especially for nonlinear PDE models, where violation of bounds can lead to instability or loss of physical accuracy. Examples include reaction-diffusion systems that model concentrations or phase-field models where maximum and minimum bounds constrain solutions.
@@
However, while DMP and monotonicity are substantial structural requirements, methods that focus on bound preservation are less restrictive and, thus, potentially more efficient for various applications. The importance of bound-preserving methods is underscored in scenarios where, e.g., positivity is required to maintain the validity of a physical model. For instance, negative solutions would be non-physical in reaction-diffusion systems or turbulence models, and possibly a stable numerical simulation can be built only by ensuring this requirement, without the need to fulfill the stronger restrictions demanded by the DMP.
@@
In the recent work [BGPV24], the idea of bound-preservation at the nodes was presented for an elliptic reaction-diffusion problem. The core idea involves choosing a baseline discretization of the PDE at hand and combining it with a projection operator that maps finite element functions to their positive (well-behaved) counterparts, ensuring that numerical solutions remain within the required bounds. The method is nonlinear, as anticipated by Godunov’s barrier theorem. For symmetric linear problems, the solution is sought as an orthogonal projection of the exact solution onto the convex set of finite element functions with positive nodal values. This work has been recently extended to convection-dominated problems in [ABP24], using continuous finite element spaces, and to time-dependent problems using discontinuous Galerkin in [BPT25]. A common feature of all the above works is the fact that they can be proven to be equivalent to solving a variational inequality over the constrained convex set (in fact, the method presented in [BPT25] is written directly in such a way, without the link to a stabilized method). This fact prevents the method from being conservative (either locally or globally), a very desirable property in, for example, time-dependent transport problems.
@@
Based on the discussion in the above paragraph, in this work we present a method that preserves the bounds of the continuous solution while at the same time is locally (and globally) conservative. The baseline discretization is the Enriched Galerkin (EG) method, in which the constraints are encoded via Lipschitz-continuous projections. The conservation property is achieved by using enriched Galerkin methods and a stabilization term that disregards element-wise constants. This approach enables the use of simple solvers, such as Richardson-like iterations, thereby avoiding more complex constrained optimization techniques commonly employed in the existing literature [EHS09, BRD+20, MN16]. Moreover, the method achieves near-best approximation properties in suitable error norms.
@@
The enriched Galerkin (EG) methods use the discontinuous Galerkin (DG) bilinear form and a combination of overall continuous and broken polynomial spaces to inherit the polynomial best approximation property from continuous finite elements and other beneficial properties, such as enhanced stability and local mass conservation from discontinuous Galerkin methods, while maintaining fewer degrees of freedom than DG. The first enriched Galerkin method was proposed by Becker et al. [BBHL03], but these methods only gained popularity after reinvention by Lee et al. [LLW16]. Since then, the EG method has been continuously generalized and applied to many problems. Its main strengths stem from its flexibility concerning the polynomial spaces that allow, for example, higher-order or spatially adaptive enrichments [RHA21, RL20]. This flexibility allows us to derive, for instance, locking-free and physics-preserving EG methods for poroelasticity [LY23, YL24] or efficient EG-based solvers for the Stokes problem [YHLA22]. However, this flexibility obstructs bound-preserving approximations even in the lowest-order cases. This issue has been addressed in [KHR20, KH23, KLY24] for hyperbolic problems and will be addressed for elliptic problems in this work.
@@
Several numerical experiments demonstrate the effectiveness of the proposed method compared to standard finite element approaches. These experiments show that the method preserves critical structural properties of the exact solution while maintaining computational efficiency. Similar methods that focus on bound preservation, such as truncation-based approaches [LHV13, Kre14, YYZ22], have been previously explored. However, the method presented here offers a fresh perspective on the formulation and numerical solution of bound-preserving problems.
The remainder of this contribution is structured as follows: Section 2 delves into the model problem and the (not bound-preserving) standard enriched Galerkin discretization. In contrast, Section 3 turns this discretization into a bound-preserving one. Section 4 is devoted to proving the existence of discrete and positive solutions, while Section 5 discusses their convergence properties to the analytic solution. Finally, Section 6 demonstrates how our method can be implemented and analyzes the condition of the linear system of equations that needs to be solved internally before a short conclusion wraps up this work.
@@ 2 Model problem and baseline discretization
We will adopt standard notation on Sobolev and Lebesgue spaces, aligned with, e.g., [EG21a]. For $D\subseteq\mathbb{R}^{d}$, $d\leq 3$, we denote by $\|\cdot\|_{0,p,D}$ the $L^{p}(D)$-norm; when $p=2$ the subscript $p$ will be omitted and we only write $\|\cdot\|_{0,D}$. In addition, for $s\geq 0$, $p\in[1,\infty]$, we denote by $\|\cdot\|_{s,p,D}$ ($|\cdot|_{s,p,D}$) the norm (seminorm) in $W^{s,p}(D)$; when $p=2$, we define $H^{s}(D)=W^{s,2}(D)$, and again omit the subscript $p$ and only write $\|\cdot\|_{s,D}$ ($|\cdot|_{s,D}$). The following space will also be used repeatedly within the text
$H^{1}_{0}(D)=\left\{v\in H^{1}(D):v=0\;{\rm on}\;\partial D\right\}.$ (1)
@@
$\int_{\Omega}\epsilon\nabla u\cdot\nabla v\;\textup{d}\boldsymbol{x}+\int_{\Omega}\mu uv\;\textup{d}\boldsymbol{x}=\int_{\Omega}fv\;\textup{d}\boldsymbol{x},$ (2)
holds for all $v\in H^{1}_{0}(\Omega)$. Using Lax-Milgram’s Lemma (see, e.g., (see, e.g., [EG21b, Lemma 25.2]), this variational problem can be shown to have a unique solution. In addition, using the comparison principle (see [RR06, Corollary 4.4]), then $u$ satisfies
$\|u\|_{0,\infty,\Omega}\leq\tilde{U}:=\max\left\{\frac{\|f\|_{0,\infty,\Omega}}{\mu},\|u_{D}\|_{0,\infty,\partial\Omega}\right\}\,.$ (3)
@@
We now describe the baseline discretization of (2). The choice made in this work is the EG method, proposed in [BBHL03, LLW16]. For this, we consider a shape-regular family of conforming simplicial triangulations $({\mathcal{T}_{h}})_{h>0}$. Only to avoid technical diversions and simplify the notations and some of the proofs, we will assume that the family of triangulations is quasi-uniform. For a triangulation ${\mathcal{T}_{h}}$, the set of its facets is denoted by ${\mathcal{F}_{h}}$. For any $T\in{\mathcal{T}_{h}}$, $h_{T}=\operatorname{diam}T$, while $h_{F}=\operatorname{diam}F$ for all $F\in{\mathcal{F}_{h}}$, $h=\max_{T\in{\mathcal{T}_{h}}}h_{T}$, and $h_{\min}=\min_{T\in{\mathcal{T}_{h}}}h_{T}$. In addition, we denote by $\boldsymbol{x}_{1},\ldots,\boldsymbol{x}_{n}$ the internal nodes of ${\mathcal{T}_{h}}$ and for $i\in\{1,\ldots,n\}$ and $T\in{\mathcal{T}_{h}}$ we define the neighborhoods
$\omega_{i}=\{T\in{\mathcal{T}_{h}}:\boldsymbol{x}_{i}\in T\}\quad\textrm{and}\quad\omega_{T}=\{T^{\prime}\in{\mathcal{T}_{h}}:T\cap T^{\prime}\not=\emptyset\}\,.$
@@
$\tilde{V}_{h}^{1}$ =$=\{v\in C(\overline{\Omega})\colon v|_{T}\in\mathbb{P}_{1}(T)\;\forall T\in{\mathcal{T}_{h}}\},$
@@
$V_{h}^{0}$ =$=\{v\in L^{2}(\Omega)\colon v|_{T}\in\mathbb{P}_{0}(T)\;\forall T\in{\mathcal{T}_{h}}\}.$
@@
$\{\!\!\{v_{h}\}\!\!\}=\tfrac{1}{2}v_{h}|_{T^{+}}+\tfrac{1}{2}v_{h}|_{T^{-}}\qquad\text{ and }\qquad[\![v_{h}]\!]=v_{h}|_{T^{+}}\boldsymbol{n}^{+}+v_{h}|_{T^{-}}\boldsymbol{n}^{-},$
@@
$\|v_{h}\|^{2}_{{\mathcal{T}_{h}}}=(v_{h},v_{h})_{\mathcal{T}_{h}}\qquad\text{ and }\qquad\|\eta_{h}\|^{2}_{{\mathcal{F}_{h}}}=\sum_{F\in{\mathcal{F}_{h}}}\frac{\epsilon+\mu h_{F}^{2}}{h_{F}}\int_{F}\eta_{h}\cdot\eta_{h}\;\textup{d}\sigma.$ (4)
Using the above notation, the standard EG method is defined as: Find $u_{h}\in V_{h}$ such that
$a_{h}(u_{h},v_{h})=b_{h}(v_{h})\qquad\text{ for all }v_{h}\in V_{h},$ (5a)
@@
$\qquad+\left\langle\gamma\frac{\epsilon+\mu h_{F}^{2}}{h_{F}^{\beta}}[\![w_{h}]\!],[\![v_{h}]\!]\right\rangle,$ (5b)
$b(v_{h})$ $=(f,v_{h})-a_{h}(I_{h}(\tilde{u}_{\textup{D}}),v_{h}).$ (5c)
Above, $\beta\in\mathbb{N}$, $\gamma=\gamma_{0}L_{\Omega}^{\beta-1}$, where $\gamma_{0}>0$ is a non-dimensional stabilization parameter, and $L_{\Omega}$ is a characteristic length of $\Omega$ (for example, it can be taken as the diameter of $\Omega$). In addition, $\tilde{u}_{\textup{D}}$ is an extension of $u_{\textup{D}}$ to the domain $\Omega$ and $\mathcal{I}_{h}:C(\overline{\Omega})\to\tilde{V}_{h}^{1}$ denotes the Lagrange interpolation operator defined in, e.g., [EG21a, Chapter 11].
Remark 1 (Properties of the EG method).
@@
    Observing that $V_{h}$ is a subspace of the standard DG space, and that $a_{h}(\cdot,\cdot)$ is an elliptic bilinear form in the DG space, and so it is elliptic in $V_{h}$ as well. More precisely, using standard arguments (see, e.g., [DPE12]), it can be shown that if $\gamma$ is large enough, there exists $c_{0}>0$, independent of $h$ and any physical constant, such that
    $a_{h}(v_{h},v_{h})\geq c_{0}\Big{(}\|\epsilon^{1/2}\nabla v_{h}\|^{2}_{0,\Omega}+\|\mu^{1/2}v_{h}\|_{0,\Omega}^{2}+\sum_{F\in{\mathcal{F}_{h}}}\gamma\frac{\epsilon+\mu h_{F}^{2}}{h_{F}^{\beta}}\|[\![v_{h}]\!]\|_{0,F}^{2}\Big{)}=:c_{0}\|v_{h}\|_{a_{h}}^{2}\,.$ (6)
@@
    The EG method is locally mass conservative, as demonstrated by testing with $v_{h}=1$ in one element $T\in{\mathcal{T}_{h}}$ and $v_{h}=0$ in all other elements $T\in{\mathcal{T}_{h}}$.
@@
    Strictly speaking, a standard symmetric interior penalty DG bilinear form would motivate to definition of the stabilization (with a varying $\beta$ scaling) by $\langle\gamma\frac{\epsilon}{h_{F}^{\beta}}[\![w_{h}]\!],[\![v_{h}]\!]\rangle.$ Thus, compared to our definition, it would not include the terms with $\mu$. While this choice will aid in the analysis, it also has a positive impact on stability with respect to the choice of $\gamma$, as discussed in the numerical section.
Remark 2.
In the above method, $\tilde{u}_{\textup{D}}$ is an extension of $u_{\mathrm{D}}$ to the interior of $\Omega$, and $I_{h}$ is the nodal Lagrange interpolation operator. Notably, we have that $u_{h}+I_{h}(\tilde{u}_{\textup{D}})\approx u$. For a standard EG discretization, the extension of $u_{\mathrm{D}}$ is irrelevant, but for the method presented below, it is simpler to consider simply an extension by zero inside the domain, and that will be the choice used herein (see [ABP24, Section 4.1] for details).

### v2

@@ Abstract
We propose a locally conservative enriched Galerkin scheme that preserves the physical bounds for an elliptic problem. To this end, we use a substantial over-penalization of the discrete solution’s jumps to obtain optimal convergence. To avoid the ill-conditioning issues that arise in over-penalized schemes, we introduce an involved splitting approach that separates the system of equations for the discontinuous solution part from the system of equations for the continuous solution part, yielding well-behaved subproblems. We prove the existence of discrete solutions and optimal error estimates, which are validated numerically.
keywords:
Maximum principle, bound preservation, local conservation, enriched Galerkin method
{MSCcodes}
65N30, 65N15, 65N12, 35J25, 35B50
@@ 1 Introduction
Physically-consistent numerical methods have become a key focus in the numerical analysis of partial differential equations (PDEs). These methods aim to preserve specific properties of the exact solutions, such as local conservation laws, entropy conditions, maximum principles, or positivity constraints. Such properties are critical for ensuring that numerical simulations of complex phenomena remain physically accurate and stable.
One particular type of physically consistent method that has received much attention are those that stay within an acceptable set of values, that is, where the discrete solution does not attain values that the continuous one would not. This is usually carried out by developing discretizations that satisfy the so-called Discrete Maximum Principle (DMP), which essentially states that the discrete solution remain within the bounds of the continuous one and does not have any local extrema that are not present in the continuous solution. This is not a new idea, and in fact in the finite element literature can be traced back to the work of Ciarlet & Raviart [CR73]. Since then, many different approaches have been proposed, and herein we refert to [Kik77, MH85, BE05, BKK08, BBK17] for some representative methods that respect the DMP in the elliptic case, and to [BJK25] for a very recent review on the topic.
@@
Now, the DMP is sometimes difficult to achieve, as in fact its validity is usually linked to mesh restrictions, and in some situations might even be too strict a requirement. In fact, sometimes a discretization that is just bound-preserving, that is, once whose solution remains within the invariant domain (i.e., the set of values that the continuous solution takes), is sufficient to guarantee stability of a simulation. In addition, methods that focus on bound preservation are less restrictive and, thus, potentially more efficient for various applications. The importance of bound-preserving methods is underscored in scenarios where, e.g., positivity is required to maintain the validity of a physical model. For instance, negative solutions would be non-physical in reaction-diffusion systems or turbulence models, and phase-field models introduce variables that need to belong to the interval [1,1][-1,1]. A stable numerical simulation can be built only by ensuring this requirement, without the need to fulfill the stronger restrictions demanded by the DMP.
@@
In the recent work [BarrenecheaGPV23], the idea of bound-preservation at the nodes was presented for an elliptic reaction-diffusion problem. The core idea involves choosing a baseline discretization of the PDE at hand and combining it with a projection operator that maps finite element functions to their positive (well-behaved) counterparts, ensuring that numerical solutions remain within the required bounds. As anticipated by Godunov’s barrier theorem, the method is nonlinear. For symmetric linear problems, the solution is sought as an orthogonal projection of the exact solution onto the convex set of finite element functions with positive nodal values. This work has been recently extended to convection-dominated problems in [AmiriBP23], using continuous finite element spaces, and to time-dependent problems using discontinuous Galerkin in [BPT25]. A common feature of all the above works is the fact that they can be proven to be equivalent to solving a variational inequality over the constrained convex set (in fact, the method presented in [BPT25] is written directly in such a way, without the link to a stabilized method). This fact prevents the method from being conservative (either locally or globally), a very desirable property in, for example, time-dependent transport problems.
@@
Based on the discussion in the above paragraph, in this work we present a method that preserves the bounds of the continuous solution while at the same time is locally (and globally) conservative. The baseline discretization is the Enriched Galerkin (EG) method, in which the constraints are encoded via Lipschitz-continuous projections. The definition of the projection is one of the main differences with the previous works [BarrenecheaGPV23, AmiriBP23, BPT25]. In fact, while in those papers the whole finite element function is restricted to belong to the acceptable interval, in the present work we split the finite element solution into its linear and constant parts. Then, the piecewise constant part remains untouched, and it is used to build the projection that will take the piecewise linear into a suitable interval. As such, the stabilization needed to remove the kernel induced by the projection vanishes in the piecewise constant functions, which guarantees local and global conservation. The method is proven to achieve near-best approximation properties in suitable error norms. In addition, this approach enables the use of simple solvers, such as Richardson-like iterations, thereby avoiding more complex constrained optimization techniques commonly employed in the existing literature [EHS09, Bochev20, MN16]. Moreover, to the authors’ knowledge, this is the first work that combines bound-preservation and local conservation with a rigorous error analysis.
@@
The enriched Galerkin (EG) methods use the discontinuous Galerkin (DG) bilinear form and a combination of overall continuous and broken polynomial spaces to inherit the polynomial best approximation property from continuous finite elements and other beneficial properties, such as enhanced stability and local mass conservation from discontinuous Galerkin methods, while maintaining fewer degrees of freedom than DG. The first enriched Galerkin method was proposed by Becker et al. [BeckerBHL03], but these methods only gained popularity after reinvention by Lee et al. [LeeLW16]. Since then, the EG method has been continuously generalized and applied to many problems. Its main strengths stem from its flexibility concerning the polynomial spaces that allow, for example, higher-order or spatially adaptive enrichments [RuppHA20, RuppL20]. This flexibility allows us to derive, for instance, locking-free and physics-preserving EG methods for poroelasticity [LeeY23, YiL24] or efficient EG-based solvers for the Stokes problem [YiHLA22]. However, this flexibility obstructs bound-preserving approximations even in the lowest-order cases. This issue has been addressed in [KuzminHR20, KuzminH23, KuzminLY24] for hyperbolic problems and will be addressed for elliptic problems in the current work.
@@
Several numerical experiments demonstrate the effectiveness of the proposed method compared to standard finite element approaches. These experiments show that the method preserves critical structural properties of the exact solution while maintaining computational efficiency. Similar methods that focus on bound preservation, such as truncation-based approaches [Lu2013TheCM, Kreuzer, YYZ22], have been previously explored. However, the method presented here offers a fresh perspective on the formulation and numerical solution of bound-preserving problems.
The remainder of this contribution is structured as follows: Section 2 delves into the model problem and the (not bound-preserving) over-penalized enriched Galerkin discretization. In contrast, Section 3 turns this discretization into a bound-preserving one. Section 4 is devoted to proving the existence of discrete and positive solutions, while Section 5 discusses their convergence properties to the analytic solution. Finally, Section 6 demonstrates how our method can be implemented and analyzes the condition of the linear system of equations that needs to be solved internally before a short conclusion wraps up this work.
@@ 2 Model problem and baseline discretization
We will adopt standard notation on Sobolev and Lebesgue spaces, aligned with, e.g., [EG21-I]. For $D\subseteq\mathbb{R}^{d}$, $d\leq 3$, we denote by $\|\cdot\|_{0,p,D}$ the $L^{p}(D)$-norm; when $p=2$ the subscript $p$ will be omitted and we only write $\|\cdot\|_{0,D}$. In addition, for $s\geq 0$, $p\in[1,\infty]$, we denote by $\|\cdot\|_{s,p,D}$ ($|\cdot|_{s,p,D}$) the norm (seminorm) in $W^{s,p}(D)$; when $p=2$, we define $H^{s}(D)=W^{s,2}(D)$, and again omit the subscript $p$ and only write $\|\cdot\|_{s,D}$ ($|\cdot|_{s,D}$). The following space will also be used repeatedly within the text
(1) $H^{1}_{0}(D)=\left\{v\in H^{1}(D):v=0\;{\rm on}\;\partial D\right\}.$
@@
(2) $\int_{\Omega}\epsilon\nabla u\cdot\nabla v\;\textup{d}\boldsymbol{x}+\int_{\Omega}\mu uv\;\textup{d}\boldsymbol{x}=\int_{\Omega}fv\;\textup{d}\boldsymbol{x},$
holds for all $v\in H^{1}_{0}(\Omega)$. Using Lax-Milgram’s Lemma (see, e.g., (see, e.g., [EG21-II, Lem. 25.2]), this variational problem can be shown to have a unique solution. In addition, using the comparison principle (see [RenardyR04, Cor. 4.4]), then $u$ satisfies
(3) $\|u\|_{0,\infty,\Omega}\leq\tilde{U}:=\max\left\{\frac{\|f\|_{0,\infty,\Omega}}{\mu},\|u_{D}\|_{0,\infty,\partial\Omega}\right\}\,.$
@@
We now describe the baseline discretization of (2). The choice made in this work is the EG method, proposed in [BeckerBHL03, LeeLW16]. For this, we consider a shape-regular family of conforming simplicial triangulations $({\mathcal{T}_{h}})_{h>0}$. Only to avoid technical diversions and simplify the notations and some of the proofs, we will assume that the family of triangulations is quasi-uniform. For a triangulation $\mathcal{T}_{h}$}, the set of its facets is denoted by $\mathcal{F}_{h}$}. For any $T\in{\mathcal{T}_{h}}$, $h_{T}=\operatorname{diam}T$, while $h_{F}=\operatorname{diam}F$ for all $F\in{\mathcal{F}_{h}}$, $h=\max_{T\in{\mathcal{T}_{h}}}h_{T}$, and $h_{\min}=\min_{T\in{\mathcal{T}_{h}}}h_{T}$. In addition, we denote by $\boldsymbol{x}_{1},\ldots,\boldsymbol{x}_{n}$ the internal nodes of $\mathcal{T}_{h}$} and for $i\in\{1,\ldots,n\}$ and $T\in{\mathcal{T}_{h}}$ we define the neighborhoods
$\omega_{i}=\{T\in{\mathcal{T}_{h}}:\boldsymbol{x}_{i}\in T\}\quad\textrm{and}\quad\omega_{T}=\{T^{\prime}\in{\mathcal{T}_{h}}:T\cap T^{\prime}\not$$\emptyset\}\,$.
@@
$\tilde{V}_{h}^{1}$ $=\{v\in C(\overline{\Omega})\colon v|_{T}\in\mathbb{P}_{1}(T)\;\forall T\in{\mathcal{T}_{h}}\},$
@@
$V_{h}^{0}$ $=\{v\in L^{2}(\Omega)\colon v|_{T}\in\mathbb{P}_{0}(T)\;\forall T\in{\mathcal{T}_{h}}\}.$
@@
[…]
@@
(4) $\|v_{h}\|^{2}_{{\mathcal{T}_{h}}}=(v_{h},v_{h})_{\mathcal{T}_{h}}\qquad\text{ and }\qquad\|\eta_{h}\|^{2}_{{\mathcal{F}_{h}}}=\sum_{F\in{\mathcal{F}_{h}}}\frac{\epsilon+\mu h_{F}^{2}}{h_{F}}\int_{F}\eta_{h}\cdot\eta_{h}\;\textup{d}\sigma.$
Using the above notation, the baseline EG method used in this work is defined as: Find $u_{h}\in V_{h}$ such that
(5a) $a_{h}(u_{h},v_{h})=b_{h}(v_{h})\qquad\text{ for all }v_{h}\in V_{h},$
@@
(5b) $\qquad+\left\langle\gamma\frac{\epsilon+\mu h_{F}^{2}}{h_{F}^{\beta}}[\![w_{h}]\!],[\![v_{h}]\!]\right\rangle,$
(5c) $b(v_{h})$ $=(f,v_{h})-a_{h}(\tilde{u}_{D,h},v_{h}).$
Above, $\tilde{u}_{D,h}\in\tilde{V}_{h}^{1}$ is an extension of the Lagrange interpolant of the boundary datum inside the domain, $\beta\in\mathbb{N}$, $\gamma=\gamma_{0}L_{\Omega}^{\beta-1}$, where $\gamma_{0}>0$ is a non-dimensional penalty parameter, and $L_{\Omega}$ is a characteristic length of $\Omega$ (for example, it can be taken as the diameter of $\Omega$).
Strictly speaking, (5) is not the standard EG formulation as our EG version enforces the element-wise constant component $u^{0}_{h}$ to approximate a homogeneous Dirichlet boundary condition (via Nitsche’s method, as enforced in DG schemes), while the overall continuous solution component $u_{h}^{1}$ is set to zero strongly on the boundary (since $V^{1}_{h}\subset H^{1}_{0}(\Omega)$). Eventually, we have that $u_{h}+\tilde{u}_{D,h}\approx u$. For a standard EG discretization, the choice of extension $\tilde{u}_{D,h}$ of the boundary datum is irrelevant, but due to its nonlinearity, for the method presented below it is simpler to follow the approach given in [AmiriBP23, Sec. 4.1] and take $\tilde{u}_{D,h}$ to vanish in all the interior nodes. This approach is also useful as we want to enforce that $u_{h}+\tilde{u}_{D,h}$ preserves the prescribed bounds at all (interior) nodes, and this is simplified by the present choice for $\tilde{u}_{D,h}$.
The aforementioned treatment of the boundary conditions allows the solution component $u^{0}_{h}$ to converge to zero rapidly, which will be a key aspect of our analysis. Thereby, $u^{0}_{h}$ is only needed to establish the local mass conservation, but does not contribute to the convergence of the method.
Remark 2.1 (Properties of the EG method).
@@
    Observing that $V_{h}$ is a subspace of the standard DG space, and that $a_{h}(\cdot,\cdot)$ is an elliptic bilinear form in the DG space, and so it is elliptic in $V_{h}$ as well. More precisely, using standard arguments (see, e.g., [PietroE12]), it can be shown that if $\gamma$ is large enough, there exists $c_{0}>0$, independent of $h$ and any physical constant, such that
    (6) $a_{h}(v_{h},v_{h})\geq c_{0}\Big(\|\epsilon^{1/2}\nabla v_{h}\|^{2}_{0,\Omega}+\|\mu^{1/2}v_{h}\|_{0,\Omega}^{2}+\sum_{F\in{\mathcal{F}_{h}}}\gamma\frac{\epsilon+\mu h_{F}^{2}}{h_{F}^{\beta}}\|[\![v_{h}]\!]\|_{0,F}^{2}\Big)=:c_{0}\|v_{h}\|_{a_{h}}^{2}\,.$
@@
    The method (5) is locally mass conservative, as demonstrated by testing with $v_{h}=1$ in one element $T\in{\mathcal{T}_{h}}$ and $v_{h}=0$ in all other elements $T\in{\mathcal{T}_{h}}$. In this case, the bilinear form $a_{h}(w_{h},1|_{T})$ is restricted to a single element $T\in{\mathcal{T}_{h}}$ and reads
    $a_{h}(w_{h},1|_{T})=\int_{T}\mu w_{h}\;\textup{d}\boldsymbol{x}+\sum_{F\in{\mathcal{F}_{h}}}\int_{F\cap\partial T}\left[\gamma\frac{\epsilon+\mu h_{F}^{2}}{h_{F}^{\beta}}[\![w_{h}]\!]-\{\!\!\{\epsilon\nabla w_{h}\}\!\!\}\right]\cdot\boldsymbol{n}_{T}\;\textup{d}\sigma.$
    The term in brackets is commonly defined as numerical flux, and it is uniquely defined on inter-element (and boundary) faces. Thus, multiplying it with the element normal $\boldsymbol{n}_{T}$ makes the face-wise integral anti-symmetric, which is the characterizing property for local mass conservation.
@@
    Strictly speaking, a standard symmetric interior penalty DG bilinear form would motivate the definition of the penalty (with a varying $\beta$ scaling) by $\langle\gamma\frac{\epsilon}{h_{F}^{\beta}}[\![w_{h}]\!],[\![v_{h}]\!]\rangle.$ Thus, compared to our definition, it would not include the terms involving $\mu$. While this choice will aid in the analysis, it also has a positive impact on stability with respect to the choice of $\gamma$, as discussed in the numerical section.